OpenStudy (anonymous):

-[5z-(14z+9)]=9+(8z+5) solve please

OpenStudy (anonymous):

-5z +14z + 9 = 8z + 5 + 9 11z + 9 = 8z +14 3z = 5 z = 5/3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

hook it up wiff dem medalz guhhh